# Modular Compilation for Quantum Chiplet Architectures

### v2

@@ 4. The Design of SEQC
We incorporate the hierarchical elements of chiplet architectures into our compiler design by adopting a two-stage procedure, as shown in Figure 2. The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
4.1. Stratification Stage
(a) A chiplet topology
@@ 4.1. Stratification Stage
In this stage, we map a quantum circuit to the inter-chiplet device topology, e.g., the structure in Figure 3(a), by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps, as shown in Figures 3(b)3(d). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
4.1.1. Qubit-to-Subcircuit Mapping
Algorithm 1 Qubit-to-Subcircuit Mapping
@@ 4.1.1. Qubit-to-Subcircuit Mapping
Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits are grouped into subcircuits, as shown in Figure 3(b). An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-chiplet SWAPs needed to transfer data between chiplets during execution. To perform this task, we leverage simulated annealing, an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989).
Simulated annealing is an iterative method to optimize a state against a given cost function. In each iteration, the state is “kicked,” or randomly perturbed, and the new state is greedily accepted if there is a reduction in cost. However, the algorithm may randomly also accept worse solutions to escape from local minima wells. The runtime of simulated annealing is statically determined using constant tuning parameters, thus the overall complexity is determined by the complexity of the cost and kick functions, which are typically polynomial. Additionally, since each run of a simulated annealing is randomized and independent, it is possible to run multiple trials in parallel with different initial random seeds, then pick the best result. This improves the chances of finding a (near-)optimal solution.
Our method is shown in Algorithm 1. We start with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined on line 1 as the number of inter-chiplet gates given a certain layout. By minimizing the number of inter-chiplet gates, we implicitly reduce the inter-chiplet connectivity. The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits. In doing so, we hope to move strongly interacting qubits close-together, minimizing the number of cross-chiplet gates. As input, our algorithm takes the input circuit, the chiplet count $N_{C}$, the chiplet size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling modifier $\beta$. The last three are tuning parameters that do not change across runs.
4.1.2. Chiplet Allocation
@@
Chiplet allocation describes the process of transforming a stratified circuit to be compatible with the inter-chiplet topology of a modular device. In Figure 2, we decompose chiplet allocation into subcircuit-to-chiplet mapping and inter-chiplet routing, analogous to how qubit allocation can be decomposed into qubit layout and routing. Accordingly, subcircuit-to-chiplet mapping, shown in Figure 3(c), defines an initial placement of logical subcircuits to physical chiplets. Subsequently, inter-chiplet routing, shown in Figure 3(d), inserts inter-chiplet SWAPs to account for limited cross-chiplet gate connectivity.
In our design, we do not explictly derive a subcircuit-to-chiplet mapping. Rather, we extend and modify the SABRE algorithm (sabre_swap) to apply to chiplets, thereby using iterations of inter-chiplet routing with randomized permutations to inform a subcircuit-to-chiplet mapping. Inter-chiplet routing seeks to consolidate the qubits of (non-SWAP) inter-chiplet gates to a single chiplet by inserting inter-chiplet SWAPs into the stratified circuit. For an single two-qubit, inter-chiplet gate, we represent inter-chiplet routing as a shortest-path problem, where the two qubits of the gate correspond to chiplet endpoints in the unweighted inter-chiplet connectivity graph. From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or chiplet distance, indicates the length of the SWAP chain.
@@
However, in the context of a entire stratified circuit, composed of multiple inter-chiplet gates, it becomes possible to interleave otherwise efficiently reorder execution of SWAP chains to reduce the depth and/or total number of inter-chiplet SWAPs. Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-chiplet routing at compile-time. Thus, we employ a similar SWAP selection scheme as SABRE (sabre_swap), where inserted SWAPs are chosen from a pool of SWAP candidates according to a cost heuristic. For our inter-chiplet routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 2. An inter-chiplet SWAP is considered symbiotic if it directly affects two inter-chiplet gates in a “beneficial” manner. In other words, each qubit of a symbiotic SWAP corresponds to a different inter-chiplet gate, where the insertion of the SWAP reduces the chiplet distance score of both gates. In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-chiplet routing cost heuristic. Next, commensalistic SWAPs are those that do not “harm,” or increase that chiplet distance, of any directly impacted gates. These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-chiplet gate without affecting other inter-chiplet gates. In certain device topologies, a SWAP acting on the qubits of two inter-chiplet gates can also exhibit commercialistic behavior. Finally, parasitic SWAPs act on the qubits of two inter-chiplet gates, where one gate benefits while the other is harmed. Since this variety of SWAP is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances to break out of a deadlock. For example, clearing one gate may free physical qubits to help route others.
4.2. Elaboration Stage
@@
(b) A $3$-qubit subcircuit post-stratification.
@@
Following the stratification stage, each chiplet can be compiled almost completely independently from one another. More precisely, the compilation process within each chiplet is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-chiplet SWAPs, which must be lowered from the chiplet level to the qubit level. Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (since each subcircuit can be processed in parallel) and problem size reduction (since there are fewer qubits to consider) of the most computationally intensive compilation stages, namely routing and optimization. To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 4 illustrates this with an example of elaboration for a single chiplet in Figure 4(a) and its corresponding subcircuit in Figure 4(b).
4.2.1. Qubit Allocation
Like other qubit allocation schemes, we incorporate layout (see Figure 4(c)) and routing (see Figure 4(e)) substages. However, the presence of inter-chiplet gates also necessitates several key changes to the SABRE (sabre_swap) qubit allocation algorithm to achieve correctness and/or maximum parallelism.
First, the placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets. To account for this, we introduce a serial stage, as shown in Figure 4(d), between the parallel layout and routing passes that greedily places gates on the nearest valid inter-chiplet edge. A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets).
@@ 4.2.1. Qubit Allocation
Once the positions of the inter-chiplet gates are settled, they must remain fixed, imposing an additional constraint on qubit routing. To promote this behavior in SABRE, we extended its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-chiplet SWAPs. Thus, we need to impose hard restrictions to ensure correctness. We establish two disjoint gates sets. One gate set operates strictly within a chiplet. This is the traditional gate set for a monolithic processor. The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs. By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
4.2.2. Parallel Translation and Optimization
@@
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity, as shown in Figure 4(f). Basis translation, in particular, is embarrassingly parallel, because every gate can be translated independently from every other gate. Meanwhile, as discussed in Section 2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
@@ 5. Methodology
To experimentally validate our compiler for modular quantum architectures, we ran a benchmark suite on a wide spectrum of circuit sizes for both SEQC and the Strawman baseline. Both SEQC and the baseline compiler were implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$. Elaboration time results were taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. When given a parallelizable workload, SEQC attempts to allocate its tasks evenly among the available physical cores, rounded to the neared whole number. Finally, the simulated annealing technique used in the qubit-to-subcircuit stage uses a starting temperature $T_{0}=200$, a rate of change of $0.005T$ each round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core.
5.1. Metrics
@@
We compare SEQC to the Strawman baseline over a wide range of common and/or relevant metrics, discussed below. We aggregate results in a manner similar to the SPEC family of benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits. The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of chiplets.
Fidelity. The top priority for quantum circuit execution is to maximize fidelity. Without a guarantee of correctness, the outputs of the quantum circuit are useless, and any speedup or quantum advantage is rendered moot. Fidelity is impacted by the decoherence constraints of individual qubits, as well as the errors on a gate-by-gate basis (mike_and_ike). Due to simulation and hardware limitations, we utilize the estimated success probability (ESP) metric (esp) to approximate fidelity. Specifically, we track the per-qubit ESP and report an aggregate ESP for each benchmark from the geometric mean of the accumulated measurements.
Non-recurring stratification time. This represents the time taken in the stratification stage, which will be present for our compiler and absent for the baseline. The stratification time represents a one-time overhead to prepare a quantum circuit for execution on a modular architecture. This result can be reused for future executions on any device with an equivalent or more expansive topology.
Recurring elaboration time. The error characteristics of a quantum device can change on a hour-by-hour basis, or even a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Thus, to generate the highest-quality circuits with the most up-to-date characteristics possible, a given quantum circuit should be recompiled for every execution. Since the qubit and chiplet topologies remain static per-device, it is possible under the SEQC framework to reuse the results from the stratification stage, and only incur overhead from repeating the elaboration stage.
Solve time. This metric is relevant only to the baseline Strawman chiplet compiler, and represents the time it takes to compile a quantum circuit. The baseline compiler solve time is a recurring cost, i.e., this compilation step has to be performed every time a circuit needs to execute on a quantum machine.
Estimated execution time. The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration. In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence. In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes.
Number of inter-chiplet gates. The inter-chiplet gates of a chiplet architecture contribute the most to error and execution time relative to other gates (at least by a factor of 4×4$\times$4 × (smith_2022_chiplets)). As a result, the number of inter-chiplet gates should represent a strong proxy metric for the quality of a compiler solution. Accordingly, minimizing the number of inter-chiplet gates should result in higher fidelity, faster-executing quantum circuits.
Circuit depth. Defined as the unweighted critical path of a quantum circuit (mike_and_ike), the circuit depth is often employed as a backend-agnostic metric for execution time that can be measured statically. Notably, it is used in SABRE (sabre_swap) as part of its cost heuristic.
Gate count. Similar to circuit depth, the total number of gates in a circuit can also be used as a static, backend-agnostic metric for certain properties of the aforementioned circuit. For instance, it may provide an indication of the amount of accumulated error, or, like circuit depth, the execution time. SABRE (sabre_swap) uses gate count to decide between randomly seeded allocation trials.
5.2. Benchmarks

### v3

@@ 3.2. The Stratify-Elaborate Quantum Compiler
We incorporate the hierarchical elements of chiplet architectures into SEQC by adopting a two-stage procedure (Figure 2). The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
3.2.1. Stratification Stage
(a) Example chiplet topology
@@ 3.2.1. Stratification Stage
In this stage, we map a quantum circuit to the inter-chiplet device topology (Figure 3(a) in our running example) by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps (Figures 3(b)3(d)). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
Qubit-to-Subcircuit Mapping
Algorithm 1 Qubit-to-Subcircuit Mapping with Sim. Annealing
@@
Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits (numbers in Figure 3(b)) are grouped into subcircuits (lowercase letters). An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-chiplet SWAPs needed to transfer data between chiplets during execution. To perform this task, we investigate two methods: Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) (ovide2024sta).
Simulated annealing is an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989). Simulated annealing is an iterative method to optimize a state against a given cost function. In each iteration, the state is “kicked,” or randomly perturbed, and the new state is greedily accepted if there is a reduction in cost. However, the algorithm may randomly also accept worse solutions to escape from local minima wells. The runtime of simulated annealing is statically determined using constant tuning parameters, thus the overall complexity is determined by the complexity of the cost and kick functions, which are typically polynomial. Additionally, since each run of a simulated annealing is randomized and independent, it is possible to run multiple trials in parallel with different initial random seeds, then pick the best result. This improves the chances of finding a (near-)optimal solution. Our method is shown in Algorithm 1. We start with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined on line 1 as the number of inter-chiplet gates given a certain layout. By minimizing the number of inter-chiplet gates, we implicitly reduce the inter-chiplet connectivity. The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits. In doing so, we hope to move strongly interacting qubits close together, minimizing the number of cross-chiplet gates. As input, our algorithm takes the input circuit $Circ$, the chiplet count $N_{C}$, the chiplet size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling factor $\beta$. The last three are constant tuning parameters across runs.
The STA algorithm is a qubit allocation algorithm designed for trapped-ion quantum systems. These devices often feature modular designs, where one device is composed of “traps” arranged in a (typically) line or ring topology, where each trap contains multiple qubits (ions). Thus, it seems possible to adapt qubit allocation algorithms designed for trapped-ion systems to support modular transmon systems as well. Specifically, one component of STA, denoted as Algorithm 2 in Ovide et al. (ovide2024sta), assigns qubits to ion traps, which we can leverage to perform qubit-to-subcircuit mapping and generate an initial subcircuit-to-chiplet mapping.
Chiplet Allocation
@@
Chiplet allocation describes the process of transforming a stratified circuit to be compatible with the inter-chiplet topology of a modular device. In Figure 2, we decompose chiplet allocation into subcircuit-to-chiplet mapping and inter-chiplet routing, analogous to how qubit allocation can be decomposed into qubit layout and routing. Accordingly, subcircuit-to-chiplet mapping (Figure 3(c)), defines an initial placement of logical subcircuits (lowercase letters) to physical chiplets (uppercase letters). Subsequently, inter-chiplet routing (Figure 3(d)) inserts inter-chiplet SWAPs to account for limited cross-chiplet gate connectivity.
In our design, we use our qubit-to-subcircuit mapping to provide an initial subcircuit-to-chiplet mapping. Then, we extend and modify the SABRE algorithm (sabre_swap) to apply to chiplets, placing iterations of inter-chiplet routing alongside randomized qubit permutations to inform a final subcircuit-to-chiplet mapping. We did explore a design that extended other components of STA to perform subcircuit-to-chiplet mapping on transmon chiplet systems, but did not produce circuits meaningfully better than this iterative approach and increased compilation time, hence it was abandoned.
@@
Inter-chiplet routing seeks to consolidate the qubits of (non-SWAP) inter-chiplet gates to a single chiplet by inserting inter-chiplet SWAPs into the stratified circuit. For a single two-qubit, inter-chiplet gate, we represent inter-chiplet routing as a shortest-path problem, where the two qubits of the gate correspond to chiplet endpoints in the unweighted inter-chiplet connectivity graph. From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or chiplet distance, indicates the length of the SWAP chain. However, in the context of a stratified circuit, composed of multiple inter-chiplet gates, it becomes possible to interleave SWAP chains (i.e., efficiently reorder their execution) to reduce the depth and/or total number of inter-chiplet SWAPs. Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-chiplet routing at compile-time. Thus, we employ a similar SWAP selection scheme as SABRE (sabre_swap), where inserted SWAPs are chosen from a pool of SWAP candidates according to a cost heuristic. For our inter-chiplet routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 2.
An inter-chiplet SWAP is considered symbiotic if it directly affects two inter-chiplet gates in a “beneficial” manner. In other words, each qubit of a symbiotic SWAP corresponds to a different inter-chiplet gate, where the insertion of the SWAP reduces the chiplet distance score of both gates. In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-chiplet routing cost heuristic. Next, commensalistic SWAPs are those that do not “harm,” or increase that chiplet distance, of any directly impacted gates. These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-chiplet gate without affecting other inter-chiplet gates. In certain device topologies, a SWAP acting on the qubits of two inter-chiplet gates can also exhibit commensalistic behavior. Finally, parasitic SWAPs act on the qubits of two inter-chiplet gates, where one gate benefits while the other is harmed. Since this variety of SWAPs is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances to break out of a deadlock. For example, clearing one gate may free physical qubits to help route others.
3.2.2. Elaboration Stage
@@
(b) A $3$-qubit subcircuit post-stratification
@@
Following the stratification stage, each chiplet can be compiled almost completely independently from one another. More precisely, the compilation process within each chiplet is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-chiplet SWAPs, which must be lowered from the chiplet level to the qubit level. Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (since each subcircuit can be processed in parallel) and problem size reduction (since there are fewer qubits to consider) of the most computationally intensive compilation stages, namely routing and optimization. To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 4 illustrates this with an example of elaboration for a single chiplet (Figure 4(a)) and its corresponding subcircuit (Figure 4(b)).
Qubit Allocation
We employ a modified version of SABRE to perform qubit allocation in the elaboration stage. Specifically, we parallelize the SABRE routing algorithm and add constraints to maintain correctness when applied to a stratified quantum circuit with inter-chiplet gates. We also explored a layout pass that used portions of the STA algorithm, but the marginal benefits to solution quality were not worth the 56×5-6$\times$5 - 6 × slower elaboration time.
@@
The placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets. To account for this, we introduce a serial stage (Figure 4(d)) between the parallel layout and routing passes that greedily places gates on the nearest valid inter-chiplet edge. A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets). Once the positions of the inter-chiplet gates are settled, they must remain fixed, imposing an additional constraint on qubit routing. To promote this behavior in SABRE, we extend its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-chiplet SWAPs. Thus, we need to impose hard restrictions to ensure correctness. We address this by establishing two disjoint gate sets. One gate set operates strictly within a chiplet. This is the traditional gate set for a monolithic processor. The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs. By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
Parallel Translation and Optimization
@@
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity (Figure 4(f)). Basis translation, in particular, is embarrassingly parallel: gates can be translated independently from each other. Meanwhile, as discussed in §2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
@@ 4. Methodology
We evaluate quantum compilers based on the quality of their solutions, their speed of compilation, and their ability to scale to large circuits/devices. To experimentally profile these characteristics, we benchmark a compiler on a suite of quantum circuits, detailed in §4.1, targeting a mock chiplet-based quantum backend, as specified in §4.2. In total, we evaluate four compilers—a reference compiler, a peephole-augmented compiler, and two versions of SEQC—on simulated backends ranging from 1 to 100 chiplets. All compilers are implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$, and timing results are taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. For our reference compiler, we use the one built into Qiskit, parameterized with optimization_level=3. Because Qiskit is not natively chiplet-aware, we run these experiments on a backend with universal inter-chiplet links (§4.2). To construct the peephole compiler we augment the reference Qiskit compiler by inserting our peephole pass between its routing and basis_translation stages, thus making it chiplet-aware. Our two SEQC versions differ in the technique they use for qubit-to-subcircuit mapping. Our first SEQC implementation leverages simulated annealing, initialized with a starting temperature $T_{0}=200$, a rate of change of $0.005T$ per round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core. Meanwhile, our second SEQC implementation leverages STA for subcircuit-to-qubit mapping and subcircuit-to-chiplet mapping.
@@
We aggregate results in a manner similar to the SPEC family of benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits. The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of chiplets.
4.1. Benchmark Suite